A 255-mL sample of nitrogen was collected over mercury at 21.0 degrees C at a pressure of 745 mm Hg. Calculate the mass of the gas.
1
Expert's answer
2020-04-27T01:45:07-0400

pV=nRT, pV= mRT/M, m= pVM/RT=745 x 255 x 28/62400 = 85.2451923.